OpenStudy (anonymous):

Limx->9 abs(9-x)/(9-x)

OpenStudy (anonymous):

\[\lim_{x\to 9}\frac{|x-9|}{x-9}\] does not exist for the following reason

OpenStudy (anonymous):

Thank you! I thought it didn't exist but wasn't 100% confident in my answer

OpenStudy (anonymous):

\[f(x) = \frac{|x-4|}{x-9} = \left\{\begin{array}{rcc} -1 & \text{if} & x <9 \\ 1 &\text{if} & x >9 \end{array} \right. \]

OpenStudy (anonymous):

oops typo there

OpenStudy (anonymous):

\[f(x) = \frac{|x-9|}{x-9} = \left\{\begin{array}{rcc} -1 & \text{if} & x <9 \\ 1 &\text{if} & x >9 \end{array} \right.\]

OpenStudy (anonymous):

so from the left the limit is \(-1\) and form the right the limit is 1